OpenStudy (anonymous):

solve the equation using square roots. 2x^2 – 128 = 0

OpenStudy (anonymous):

2x^2 + 39 = 57 this one too

OpenStudy (anonymous):

2x^2 – 128 = 0 2*x^2 = 128 x^2 = 128/2 x^2 = 64 x = sqrt(64) x = +64 or -64

OpenStudy (anonymous):

2x^2=128 x^2=64 x=+-8

OpenStudy (anonymous):

yes, my bad. i forgot to square root 64.